### v5

# Foundation Models on a Budget: Approximating Blocks in Large Vision Models

Abstract
Foundation Models have shown impressive performance in various tasks and domains, yet they require massive computational resources, raising concerns about accessibility and sustainability. Previous attempts to reduce foundation model size fall short of fully addressing the problem, as they end up increasing computational load through additional training steps. Recent works reveal that deep neural networks exhibit internal representation similarities. While inter-network similarities have enabled techniques such as model stitching and merging, intra-network similarities remain underexplored for improving efficiency. In this paper, we propose Transformer Blocks Approximation (TBA), a novel method that leverages intra-network similarities to identify and approximate transformer blocks in large vision models. TBA replaces these blocks using lightweight, closed-form transformations, without retraining or fine-tuning the rest of the model. The proposed method reduces the number of parameters while having minimal impact on the downstream task. We validate the effectiveness and generalizability of TBA through extensive experiments across multiple datasets (e.g., Imagenet-1k and CIFAR100) and state-of-the-art pretrained vision models (e.g, ViT, DiNO-v2, and DEiT).
1 Introduction
As Neural Networks (NNs) grow in size and complexity, their demand for computational resources has become a significant bottleneck. Despite the impressive performance of large models, they often come with substantial trade-offs, such as increased memory and power consumption. This has led to a growing interest in methods that can reduce model complexity without sacrificing performance. However, most existing approaches trying to mitigate these challenges either demand additional training steps or result in substantial performance degradation. Recent research showed that there exist similarities between representation within and between NNs. The observation that many layers or components within these networks perform similar functions or produce highly correlated outputs suggests that some blocks may be approximated, highlighting opportunities for network simplification. In this paper, we investigate two key research questions: (i) how to effectively identify blocks within neural networks that yield similar representations, and (ii) how to approximate these blocks with minimal impact on overall performance without any additional training steps. To address the first question, we provide an extensive analysis showing that the Mean Squared Error (MSE) effectively identifies such blocks, suggesting that their approximation should have a negligible impact. For the second question, we propose Transformer Blocks Approximation (TBA), a novel method that leverages similarities between representations to approximate transformer blocks using lightweight transformations, such as linear mappings. Once the blocks that have minimal impact on model functionality are identified, TBA replaces them with a simpler transformation. This approximation strategy allows for reducing the overall parameter count, ensuring minimal impact on downstream task performance. Our main contributions are as follows:
We introduce TBA, a general framework for identifying and approximating transformer blocks that produce similar outputs, using simpler transformations (e.g., linear), reducing model parameters and complexity with minimal to no impact on the downstream task performance (Figure 1).
2 Related work
Measuring similarities A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023). One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations. Variants of CCA, such as Singular Value CCA (SVCCA) (Raghu et al., 2017), aim to enhance robustness, while techniques like Projection Weighted CCA (PWCCA) (Morcos et al., 2018) mitigate sensitivity to small perturbations. Another widely used metric, Centered Kernel Alignment (CKA) (Kornblith et al., 2019), captures the similarity between latent spaces while ignoring orthogonal transformations. However, recent work (Davari et al., 2022) highlights that this metric can be sensitive to shifts in the latent space. Additionally, Barannikov et al. (2021) proposes a method to compare two data representations by measuring the multi-scale topological dissimilarity, while Fumero et al. (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
Leveraging similarities Valeriani et al. (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models. Kvinge et al. (2022) investigate how models process variations in data points across layers, while Nguyen et al. (2020) assess the impact of network depth and width on hidden representations. Additionally, Crisostomi et al. (2023) study the conditions under which two latent spaces can be merged into a unified one. Moschella et al. (2023) construct a unified space shared by different NNs, enabling zero-shot stitching of independently trained models across different modalities (Norelli et al., 2023). More recently, Cannistraci et al. (2024) enable model stitching without explicit assumptions about the transformation class connecting the latent manifold embeddings, or with only partial correspondence between latent spaces (Cannistraci et al., 2023). Finally, Lähner & Moeller (2024); Maiorca et al. (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
Architectural efficiency While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs. Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research. For instance, Veit et al. (2016) shows that removing residual blocks from deep Convolutional Neural Networks (CNNs) only marginally impacts performance, which inspired approaches to reduce inference time by dynamically deciding which layers to execute based on the input (Wu et al., 2018; Veit & Belongie, 2018). Additionally, various techniques to enhance efficiency have emerged, such as early exiting and model pruning. Early exit strategies, which introduce intermediate output layers at different stages of the network, have been shown to reduce inference time (Xin et al., 2020; Zhou et al., 2020; Yu et al., 2022; Tang et al., 2023). However, these approaches require the training of intermediate classifiers to enable exits at predefined layers. Alternatively, model pruning reduces computational load by either removing individual weights based on specific criteria, such as gradient information (Ma et al., 2023), entropy (Liao et al., 2023), or second-order information (Singh & Alistarh, 2020), or by eliminating larger structural components, like channels or residual blocks in ResNets (Bai et al., 2023; Wang & Wu, 2023) and self-attention layers in Transformers (Zhang & He, 2020; Sajjad et al., 2023; Venkataramanan et al., 2024; Zhang et al., 2024). Although effective, these approaches require training the model from scratch and, in the best case, fine-tuning. However, Bai et al. (2023) shows that for CNNs, this additional training step can sometimes be avoided.
Unlike other methods, TBA leverages intra-network similarities to reduce foundation model complexity without the need for additional training steps while maintaining competitive performance.
3 Transformer Blocks Approximation
The core idea of the proposed approach is to detect similar representations within NNs and replace the corresponding blocks with simple transformations. In the context of this work, a "block" refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, functioning as a cohesive unit. Instead of executing the entire NN, using transformer block approximation (TBA), we approximate these blocks, reducing computational complexity while preserving the network’s functionality. To provide a clear understanding, Figure 1 presents an overview of our method.
Identifying similar representations We hypothesize that certain foundation models may contain blocks that produce similar representations. To quantify these similarities, we employ the MSE. A low MSE between the output of an earlier block and a later block indicates that their respective representations are highly similar. This suggests that the intervening sequence of blocks contributes minimally to transforming the representation, highlighting potential redundancy.
Let represent the total number of blocks in the model. Let denote the output representation (a vector of dimension ) of block for a single input, where . For any two blocks, a starting block and an ending block (where ), we compute their output representations and . For each , where is a subset of points sampled uniformly at random from , defines the input to the corresponding layer . The MSE between these representations is defined as:
By systematically evaluating the MSE for various pairs of blocks , we can identify those being good candidates for approximation. This allows for targeted complexity reduction without substantially altering the network’s internal representations or downstream-task performance at a very low computational cost.
Approximating transformer blocks Once two blocks and are identified as having highly similar output representations, our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block . This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs. This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with .
Let and represent the output representations from block and respectively, for the data points in . Our objective is to find a transformation such that:
In this work, we consider to be a linear transformation that can be estimated by minimizing the squared error between the transformed output and the actual , which can be solved using least squares:
This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation . The solution bypasses the computation of all layers between any two blocks and , replacing them with . This approximation significantly reduces computational complexity by replacing one or more transformer blocks, comprising multi-head self-attention and feed-forward layers, with a low-cost linear transformation. To sum up, our approach consists of two main stages:
1.
Identification: We use MSE to identify blocks that produce similar representations.
2.
Approximation: We compute an efficient linear approximation using a transformation matrix to bypass the previously identified blocks.
This procedure reduces parameter count with minimal impact on internal representations, as shown in Figures 9, 7, 10, 11 and 8, and maintains compatibility with any downstream classifiers, achieving significant compression as shown Tables 1, 8, 9, 10, 11, 12 and 13.
4 Experiments
This section begins with an analysis of block-level similarities within pre-trained vision foundation models (Section 4.1), and we investigate how TBA impacts the representations of these latent spaces. Following this, Section 4.2 provides a comprehensive quantitative evaluation of TBA’s performance on various image classification tasks, including standard, zero-shot, and cross-dataset generalization scenarios. Lastly, in Section 4.3 we compare TBA against existing baselines, including an ablation study on transformation complexity, further validating the effectiveness of the proposed approach.
4.1 Latent analysis
Block similarities We investigate similarities emerging in latent representations using three distinct transformer-based models with different dimensionality: ViT-S, DiNO-B, and DEiT-S, and five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k.
Details on the models and datasets are provided in Table 6 and Table 7. Figure 2 presents the MSE matrices between blocks of different models when using different datasets. These similarities are calculated using only a small subset of the training data (i.e., 500 samples). Results reveal that while the patterns of similarity vary across models, they remain consistent across different datasets. This suggests that the similarity structure between computational blocks is predominantly influenced by the model rather than the dataset used. This finding aligns with observations from Nguyen et al. (2020), where DNNs trained from scratch tend to exhibit a distinctive “block structure” in their representations, linked to model overparameterization. Our results extend this observation to pretrained foundation models, where we show that such a structure is primarily influenced by the model. We conduct the analysis using the mean over the tokens, but in Section A.2.1 we show that this also applies when using only the CLS token from each block. Finally, in Figure 6, we present an ablation study to investigate the influence of the similarity metric. We correlate various metrics with downstream task performance using ViT-S and ImageNet1k. Our findings reveal that using simple metrics such as MSE or cosine similarity to identify block-level similarities exhibits a strong correlation with final accuracy. These results underscore the flexibility of our proposed method, indicating its compatibility with a range of similarity or distance functions that can capture inter-block relationships.
Effects of TBA on latent representations After identifying blocks producing similar representations, we analyze the impact on the latent representation output by the last block when approximating using TBA.
These blocks are approximated using a shared linear transformation applied across all tokens based on a subset of 3,000 training samples. For consistency, we employed the same models and datasets used for the results in Figure 2. We compute the MSE between the representations in the last block of the original and the TBA model when approximating the block. As illustrated in Figure 3, in most cases, the MSE increases as the block depth increases.
Original TBA
DiNO-B
DEiT-S
This suggests that approximating the final blocks would lead to significant changes in the final representations, indicating their critical role. However, in the case of DEiT-S, the trend is different. Here, the MSE decreases toward the final blocks. This is also shown in the similarity matrices illustrated in Figure 2, where the final blocks have a high similarity. These findings reinforce the intuition behind the use of MSE, demonstrating a correlation between the metric and the final representation similarity when approximating blocks. In some instances, such as with the MNIST dataset, the MSE scores remain relatively consistent across blocks, indicating that the representations are largely similar to one another. However, for more complex datasets like ImageNet1k, the representations in the final or the first blocks become increasingly dissimilar, making it advantageous to approximate intermediate blocks. This suggests that the similarities are primarily driven by the model but also guided by the complexity of the dataset, allowing for targeted approximations that reduce model parameters and complexity without significantly compromising performance. To further investigate the effect of TBA on the latent representations, we visualize, using Principal Component Analysis (PCA), the representations of the TBA model when approximating the final block and its original representation (see Figure 4). The plot shows the representations generated using the DiNO-B and DEiT-S pretrained encoders on F-MNIST. Colors represent classes. The visualization highlights that approximating the final block of DiNO-B results in noticeable deviations from the original representations, while in DEiT-S the approximated representations remain similar to the original ones. This observation aligns with the results from Figure 3, where approximating blocks can lead to significant changes in representations. For complete results, refer to Section A.2.2.
4.2 Transformer block approximation performance
We perform image classification using various pretrained models with different dimensionalities (ViT-T, ViT-S, ViT-L, ViT-B, DiNO-S, DiNO-B, and DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C, CIFAR-100F, and ImageNet1k), keeping all models frozen. Block approximations are computed using a shared linear transformation applied across all tokens, based on
| Accuracy | |||||
| Approx. | Params. | CIFAR-10 | CIFAR-100F | ImageNet1k | |
| ViT-S | 1 5 | -6.51M | (-11.09) | (-27.35) | (-39.86) |
| 2 5 | -4.88M | (-4.88) | (-14.42) | (-17.56) | |
| 7 10 | -4.88M | (-8.69) | (-23.89) | (-51.14) | |
| 2 4 | -3.26M | (-2.95) | (-8.67) | (-6.80) | |
| 9 11 | -3.26M | (-5.00) | (-14.71) | (-35.54) | |
| 1 2, 4 5 | -3.26M | (-3.54) | (-11.10) | (-12.06) | |
| 0 1 | -1.63M | (-2.09) | (-6.15) | (-11.10) | |
| 3 4 | -1.63M | (-0.66) | (-2.47) | (-2.51) | |
| 9 10 | -1.63M | (-1.15) | (-6.19) | (-15.58) | |
| 10 11 | -1.63M | (-1.67) | (-4.97) | (-12.73) | |
| original | 21.82M | ||||
| DiNO-B | 1 5 | -26.5M | (-11.28) | (-28.62) | (-58.57) |
| 2 5 | -19.5M | (-5.87) | (-19.50) | (-28.26) | |
| 7 10 | -19.5M | (-4.46) | (-14.02) | (-35.88) | |
| 2 4 | -13M | (-1.49) | (-6.62) | (-7.70) | |
| 9 11 | -13M | (-10.76) | (-14.36) | (-27.74) | |
| 1 2, 4 5 | -13M | (-1.47) | (-6.11) | (-7.79) | |
| 0 1 | -6.5M | (-0.05) | (-1.01) | (-1.00) | |
| 3 4 | -6.5M | (-0.38) | (-2.78) | (-1.62) | |
| 9 10 | -6.5M | (-0.78) | (-3.26) | (-10.09) | |
| 10 11 | -6.5M | (-1.01) | (-4.23) | (-13.83) | |
| original | 86.58M | ||||
| DEiT-S | 1 5 | -6.51M | (-13.55) | (-28.90) | (-40.88) |
| 2 5 | -4.88M | (-5.41) | (-14.87) | (-16.05) | |
| 7 10 | -4.88M | (-1.61) | (-2.78) | (-22.24) | |
| 2 4 | -3.26M | (-1.85) | (-6.37) | (-5.32) | |
| 9 11 | -3.26M | (+0.30) | (+1.11) | (-5.39) | |
| 1 2, 4 5 | -3.26M | (-5.74) | (-13.31) | (-10.68) | |
| 0 1 | -1.63M | (-6.21) | (-12.91) | (-15.38) | |
| 3 4 | -1.63M | (-0.14) | (-1.24) | (-1.22) | |
| 9 10 | -1.63M | (+0.30) | (+0.86) | (-2.15) | |
| 10 11 | -1.63M | (+0.49) | (+1.15) | (+0.05) | |
| original | 21.82M | ||||
a subset of 3,000 training samples. Then, we train a single linear layer using the Adam optimizer (learning rate ) for 5 epochs, with 3 seeds and a batch size of 256. Details regarding the models, datasets, computational resources, and software tools are provided in Table 6, Table 7, Section A.1.4, and Section A.1.3, respectively.
Image classification As presented in Table 1, the proposed TBA method successfully reduces model size while maintaining, and in some cases even improving, image classification performance. These downstream task results support the findings from the analysis in Section 4.1. Specifically, as discussed in Section 4.1 and illustrated in Figures 3 and 4, approximating the final block of DEiT-S yields a final representation highly similar to the original. This similarity suggests such blocks are optimal candidates for approximation, a conclusion further validated by Table 1. The table shows that even when approximating multiple consecutive blocks (e.g., 911), we can reduce parameters while achieving performance superior to the original model. Overall, the consistent or enhanced performance indicates that a linear transformation is sufficient to approximate these transformer blocks, thereby effectively reducing model parameters. It’s important to note that this transformation is shared across all tokens, with no additional training steps required. Additional results are presented in Tables 8, 9, 10, 11, 12 and 13. To assess the scalability of our method, we extended our evaluation to ViT-L. While our approach reduces its parameters, ViT-L still imposes considerable computational demands, necessitating high-performance GPU infrastructure. This finding underscores that while effective for substantial compression, our method’s impact on making extremely large models practical for resource-constrained environments is limited by their initial scale. Finally, in Section A.2.6, we further analyze model behavior after block approximations on the final image-classification task, complementing the quantitative results with qualitative insights into the approximation procedure.
Zero-shot image classification To further assess the effectiveness of our approach, we evaluate
| Params. | Approx. | Accuracy | |
| -6.49M | 0 1 | 57.93 | - |
| 1 2 | 64.20 | - | |
| 2 3 | 66.35 | -5.51% | |
| 3 4 | 64.65 | - | |
| 4 5 | 64.86 | - | |
| 5 6 | 58.05 | - | |
| 6 7 | 61.56 | - | |
| 7 8 | 58.53 | - | |
| 8 9 | 52.32 | - | |
| 9 10 | 59.21 | - | |
| 10 11 | 22.64 | - | |
| 149.07M | original | 70.21 | – |
TBA in a zero-shot image classification setting. This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al. (2022), with ImageNet1k serving as the downstream evaluation dataset. As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples. Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged. We follow the standard ImageNet1k prompt templates. As shown in Table 2, when approximating earlier vision blocks (e.g., 1 2 or 2 3), the model still achieves competitive results, but slightly reduces the number of parameters. The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper. Details on the models are provided in Table 6.
Transformation generalization We further evaluate the model’s ability to generalize transformations across datasets. Specifically, we analyze the capability of applying a linear transformation learned on one dataset (source) to another (target), using the same underlying pretrained architecture and without any subsequent training or fine-tuning of the approximated model. Table 3 reports accuracy scores for ViT-S and DiNO-S on MNIST, CIFAR-10, and CIFAR-100, where the "Source" column indicates the source dataset used to compute the transformation. The results demonstrate that a single, token-shared linear transformation can generalize effectively across different target datasets, with the exception of MNIST, which may be too simplistic to yield broadly transferable representations. This cross-dataset generalization is particularly valuable in scenarios where fine-tuning is infeasible, such as in privacy-sensitive domains like healthcare or finance. In these contexts, regulatory constraints often restrict training on local data, yet leveraging existing pretrained models remains crucial. Our findings suggest that lightweight adaptation via shared linear transformations offers a promising path for improving model performance on custom datasets without requiring additional training. Additional results are provided in Tables 14 and 15.
| Accuracy | ||||
| Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C |
| 3 4 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
| 3 5 | MNIST | |||
| CIFAR-10 | ||||
| CIFAR-100F | ||||
4.3 Comparison
Direct comparisons for TBA are limited, as existing approaches either fine-tune pretrained models or train from scratch. In contrast, TBA maintains the foundation model frozen. We include two types of comparisons: a variant of SkipAt, a method proposed in Venkataramanan et al. (2024), and an ablation study on transformation complexity.
SkipAt identity Venkataramanan et al. (2024) explores the effect of skipping self-attention or multi-head self-attention by inserting identity functions in pretrained models, without retraining. We extend this idea by skipping entire blocks, not just attention components. As shown in Table 4, TBA consistently outperforms this SkipAt Identity baseline on CIFAR-10 and CIFAR-100F. The accuracy improvements confirm that approximating is more effective than skipping them. Results with other datasets and models can be found in Section A.2.5.
| CIFAR-10 | CIFAR-100F | |||
| Approx. | SkipAt | TBA | SkipAt | TBA |
| 1 5 | ||||
| 2 5 | ||||
| 7 10 | ||||
| 1 3 | ||||
| 2 4 | ||||
| 9 11 | ||||
| 2 3 | ||||
| 9 10 | ||||
| original | ||||
Trained approximators We also evaluate whether more complex approximations improve performance. Specifically, we compare TBA to deeper MultiLayer Perceptron (MLP)-based translators (MLP and Res-MLP), which are trained for 300 steps using Adam with a learning rate of 1e-3. As shown in Table 5, employing a simple linear transformation to approximate blocks is the optimal choice in most cases. While deeper translators occasionally yield slightly higher accuracy (e.g., 15), they incur significant training overhead and require gradient-based optimization. In contrast, TBA is entirely training-free and operates in closed form; no backpropagation or updates are required. This highlights the efficiency–performance trade-off and supports the design choice behind TBA. Refer to Section A.1.2 for approximator details.
| Approx. | TBA | MLP | Res-MLP |
| 1 5 | |||
| 2 5 | |||
| 7 10 | |||
| 1 3 | |||
| 2 4 | |||
| 9 11 | |||
| 3 4 | |||
| 9 10 |
5 Limitations and future work
While TBA demonstrates significant promise in efficiently approximating transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks. Future research will explore the applicability of TBA to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction). Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TBA’s adaptability. Furthermore, we aim to expand the analysis of these block-level similarities. This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward network layers, and potentially at coarser granularities across larger segments of the network. Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus last token). The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.
6 Conclusion
In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller yet performant models. To this end, we propose Transformer Blocks Approximation (TBA), a novel method for identifying and efficiently approximating similar transformer blocks using a simple linear transformation, without requiring additional training or fine-tuning. Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TBA significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance. Notably, the learned linear transformations exhibit strong generalization capabilities across datasets, enabling lightweight model adaptation. Furthermore, TBA’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations. TBA thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible.
Acknowledgments
The authors gratefully acknowledge Luca Moschella for the insightful discussions. TS and EP are supported by the grant #2021-911 of the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). EP is also supported by a fellowship from the ETH AI Center. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI).
Appendix A Appendix
A.1 Implementation details
This section details the experiments conducted in Section 4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
A.1.1 Models and Datasets
Table 6 contains the full list of the pretrained models, while Table 7 contains dataset information.
| Modality | HuggingFace Model Name | Alias | Enc. Dim |
| Vision | WinKawaks/vit-tiny-patch16-224 | ViT-T (Dosovitskiy et al., 2021) | 192 |
| WinKawaks/vit-small-patch16-224 | ViT-S (Dosovitskiy et al., 2021) | 384 | |
| facebook/dinov2-small | DiNO-S (Oquab et al., 2023) | 384 | |
| facebook/deit-small-patch16-224 | DEiT-S (Touvron et al., 2020) | 384 | |
| google/vit-base-patch16-224 | ViT-B (Dosovitskiy et al., 2021) | 768 | |
| facebook/dinov2-base | DiNO-B (Oquab et al., 2023) | 768 | |
| laion/CLIP-ViT-B-16-laion2B-s34B-b88K | OpenCLIP-ViT-B (Zhai et al., 2019) | 768 | |
| google/vit-large-patch16-224 | ViT-L (Dosovitskiy et al., 2021) | 1024 |
| Modality | Name | Alias | Classes |
| Vision | MNIST (Deng, 2012) | MNIST | 10 |
| Fashion-MNIST (Xiao et al., 2017) | F-MNIST | 10 | |
| CIFAR-10 (Krizhevsky et al., 2009) | CIFAR-10 | 10 | |
| CIFAR-100 (coarse) (Krizhevsky et al., 2009) | CIFAR-100C | 20 | |
| CIFAR-100 (fine) (Krizhevsky et al., 2009) | CIFAR-100F | 100 | |
| Imagenet-1k (Russakovsky et al., 2015) | ImageNet1k | 1000 |
A.1.2 Approximators
The first implementation, referred to as the Res-MLP, is composed of two normalization layers and a feedforward submodule. The first layer normalization processes the input tensor, followed by a feedforward submodule comprising a linear transformation, a SiLU activation, a dropout layer, and a final linear transformation. The output of the feedforward submodule is added to the normalized input via a residual connection. This sum is then passed through the second normalization layer to produce the final output. The second implementation, referred to as the MLP, is a simplified MLP that employs a sequential architecture with a first linear transformation that reduces the input dimensionality to half of the target dimension, followed by a GELU activation function for smooth non-linearity, and a final linear transformation that restores the reduced features to match the target dimensionality. Refer to LABEL:lst:mlp1 and LABEL:lst:mlp2 for the code snippet of the two translators.
A.1.3 Tools & technologies
All the experiments presented in this work employ the following tools:
PyTorch Lightning, to ensure reproducible results while also getting a clean and modular codebase;
NN-Template GrokAI (2021), to easily bootstrap the project and enforce best practices;
Transformers by HuggingFace, to get ready-to-use transformers for both text and images;
Datasets by HuggingFace, to access most of the datasets;
DVC (Kuprieiev et al., 2022), for data versioning;
A.1.4 Computational resources
Experiments involving larger models, specifically DiNO-B, OpenCLIP-ViT-B, and ViT-L, were conducted on an NVIDIA H100 GPU equipped with 93 GB of memory. All the other experiments utilized an NVIDIA GeForce RTX 5090 GPU with 31 GB of memory.
A.2 Additional Experiments
This section presents supplementary experiments to extend those detailed in Section 4.
A.2.1 Block Similarities
The analysis in Section 4.1 reveals distinct block-wise similarity patterns within pretrained foundation models. The analysis utilizes the mean of token representations within each block to compute similarities. However, as detailed in Figure 5, these characteristic similarity patterns persist and are qualitatively very similar when using only the CLS token’s representation from each block. This suggests a degree of interchangeability for this type of analysis.
Finally, to assess the impact of the similarity measure itself, we conducted an ablation study correlating various metrics with downstream task performance on ViT-S and ImageNet1k (Figure 6). Our findings reveal that even simple metrics, such as MSE or cosine similarity, when used to identify block-level similarities, exhibit a strong correlation with final accuracy. These results underscore the robustness and flexibility of leveraging such block similarities, indicating compatibility with a range of similarity or distance functions capable of capturing these inter-block relationships.
| Approx. | Params. | ImageNet1k |
| 1 5 | 15.31M | |
| 2 5 | 16.94M | |
| 7 10 | 16.94M | |
| 1 3 | 18.56M | |
| 3 5 | 18.56M | |
| 2 4 | 18.56M | |
| 9 11 | 18.56M | |
| 2 3 | 20.19M | |
| 3 4 | 20.19M | |
| 4 5 | 20.19M | |
| 9 10 | 20.19M | |
| - | 21.82M |
A.2.2 PCA Visualization
In Section 4.1 we show the effect of TBA on latent representations. In this section, we provide additional visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures 10, 11, 9, 7 and 8).
A.2.3 Image Classification
This section presents additional experiments that complement and extend those detailed in Section 4.2. Datasets and models are the ones detailed in Tables 7 and 6.
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Accuracy | ||||||
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | -25.99M | |||||
| 2 5 | -19.49M | |||||
| 1 3 | -13M | |||||
| 3 5 | -13M | |||||
| 2 4 | -13M | |||||
| 8 10 | -13M | |||||
| 9 11 | -13M | |||||
| 3 4 | -6.5M | |||||
| 4 5 | 6.5M | |||||
| 0 1 | -6.5M | |||||
| 1 2 | -6.5M | |||||
| 2 3 | -6.5M | |||||
| - | 86.39M | |||||
| Accuracy | |||||||
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 112 | -127.02M | ||||||
| 04 | -46.19M | ||||||
| 15 | -46.19M | ||||||
| 25 | -34.64M | ||||||
| 710 | -34.64M | ||||||
| 02, 34 | -34.64M | ||||||
| 12, 34 | -23.09M | ||||||
| 13 | -23.09M | ||||||
| 24 | -23.09M | ||||||
| 35 | -23.09M | ||||||
| 57 | -23.09M | ||||||
| 01 | -11.55M | ||||||
| 12 | -11.55M | ||||||
| 23 | -11.55M | ||||||
| 34 | -11.55M | ||||||
| 45 | -11.55M | ||||||
| 56 | -11.55M | ||||||
| 67 | -11.55M | ||||||
| 78 | -11.55M | ||||||
| 89 | -11.55M | ||||||
| 910 | -11.55M | ||||||
| 1011 | -11.55M | ||||||
| 1112 | -11.55M | ||||||
| 1213 | -11.55M | ||||||
| 1314 | -11.55M | ||||||
| 1415 | -11.55M | ||||||
| 1516 | -11.55M | ||||||
| 1617 | -11.55M | ||||||
| 1718 | -11.55M | ||||||
| 1819 | -11.55M | ||||||
| 1920 | -11.55M | ||||||
| 2021 | -11.55M | ||||||
| 2122 | -11.55M | ||||||
| 2223 | -11.55M | ||||||
| original | 304.35M | ||||||
A.2.4 Transformation Generalization
In this section, we extend the analysis to evaluate the model’s ability to generalize transformations across datasets to different approximation blocks and models (i.e., DiNO-S).
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| ViT-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| Accuracy | ||||||
| Encoder | Approx. | Source | MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| DiNO-S | 2 3 | MNIST | ||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 4 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 3 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 3 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 4 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 8 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 9 11 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 2 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 7 10 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
| 1 5 | MNIST | |||||
| CIFAR-10 | ||||||
| CIFAR-100 | ||||||
A.2.5 Comparison
In this section, we provide additional results for the SkipAt baseline.
| Accuracy | |||||
| Skip | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | |||||
| 2 5 | |||||
| 7 10 | |||||
| 1 3 | |||||
| 3 5 | |||||
| 2 4 | |||||
| 8 10 | |||||
| 9 11 | |||||
| 0 1 | |||||
| 1 2 | |||||
| 2 3 | |||||
| 3 4 | |||||
| 4 5 | |||||
| 5 6 | |||||
| 6 7 | |||||
| 7 8 | |||||
| 8 9 | |||||
| 9 10 | |||||
| 10 11 | |||||
| original | |||||
A.2.6 Analysis of misclassifications
In this section, we examine changes in per-class accuracy and misclassification patterns. As shown in Figure 12, models behave differently at block approximations. DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation). ViT-S shows a similar drop for class dog on its final block. Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated. For DEiT-S, several mid-to-late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class. These observations suggest strategies like preferring late-block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation. In Figure 13, we show the delta confusion matrix for DEiT-S on CIFAR-100C. Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal). On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal). This supports results shown in Figures 12 and 1.
Additionally, Figure 14 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S. The patterns we observe mirror the trends in Figures 12 and 13: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified. Instead, when approximating later blocks, we observe images of the class dog to be misclassified. Together, these qualitative examples show that understanding these block-specific vulnerabilities allows us to steer the approximation procedure, informing choices about which blocks to approximate based on the observed impact on the final model’s class-wise performance.