### v6

# : Transformer Optimization using Adaptive and Simple Transformations

Abstract
Foundation models achieve State-of-the-art (SOTA) performance across different tasks, but their size and computational demands raise concerns about accessibility and sustainability. Existing efficiency methods often require additional retraining or fine-tuning, limiting their practicality. Recent findings suggest that deep neural networks exhibit internal representation similarities. While such similarities across different models have been exploited for enabling techniques such as model stitching and merging, intra-network redundancy remains underexplored as a source for efficiency gains. In this paper, we introduce Transformer Optimization using Adaptive and Simple Transformations (TOAST), a framework that exploits these redundancies to approximate entire transformer blocks with lightweight closed-form mappings, such as linear transformation or even the identity, without any additional training. Across SOTA pretrained vision models (e.g., ViT, DINOv2, DeiT) and datasets ranging from MNIST to ImageNet-1k, TOAST reduces parameters and computation while preserving, and in some cases improving, downstream performance. These results show that large portions of transformer depth can be replaced by trivial functions, opening a new perspective on efficient foundation models.
1 Introduction
As Neural Networks (NNs) continue to grow in size and complexity, their demand for computational resources has become a critical bottleneck. While larger models consistently achieve SOTA performance, this comes at the cost of substantial memory usage and power consumption, limiting their accessibility and deployment. This challenge is for instance most relevant in on-device scenarios, where saving memory, latency, and energy, even by little margins, is critical (Pan et al., 2022; Li et al., 2022). This has motivated a growing body of work on reducing model complexity. However, most existing approaches either require additional, resource-intensive training phases or lead to significant drops in accuracy. Recent studies reveal that there exists strong representational similarities both within and between NNs. In other words, when focusing on intra-network similarities, different blocks often perform overlapping functions or produce highly correlated outputs.
This redundancy suggests an opportunity: instead of retraining or pruning, can we approximate these blocks with simpler transformations? To address this question, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel framework that exploits block-level representational redundancy to replace transformer blocks with lightweight transformations. By doing so, TOAST reduces parameter count and computational cost, while maintaining (and in some cases even improving) downstream task performance. Crucially, our method is training-free, making it simple, efficient, and widely applicable, even in resource-constrained scenarios such as deployment on edge devices, where even the smallest available models may exceed memory or power budgets. Our main contributions are as follows:
We propose TOAST, a simple yet effective framework that replaces transformer blocks with lightweight transformations (e.g., linear maps or even the identity), significantly reducing parameters and computational cost while preserving downstream performance (Figure˜1).
2 Related work
Measuring Similarities
A range of metrics have been introduced to assess the similarity between latent spaces generated by different NNs (Klabunde et al., 2023; Ballester et al., 2023). One established approach is Canonical Correlation Analysis (CCA) (Hotelling, 1992), known for its invariance to linear transformations. Variants of CCA, such as Singular Value CCA (SVCCA) (Raghu et al., 2017), aim to enhance robustness, while techniques like Projection Weighted CCA (PWCCA) (Morcos et al., 2018) mitigate sensitivity to small perturbations. Another widely used metric, Centered Kernel Alignment (CKA) (Kornblith et al., 2019), captures the similarity between latent spaces while ignoring orthogonal transformations. However, recent work (Davari et al., 2022) highlights that this metric can be sensitive to shifts in the latent space. Additionally, Barannikov et al. (2021) proposes a method to compare two data representations by measuring the multi-scale topological dissimilarity, while Fumero et al. (2024) leverages the principles of spectral geometry to model and analyze the relationships between distinct latent spaces.
Leveraging Similarities
Valeriani et al. (2024) examine the intrinsic dimensionality and neighbor compositions of representations in transformer models. Kvinge et al. (2022) investigate how models process variations in data points across layers, while Nguyen et al. (2020) assess the impact of network depth and width on hidden representations. Additionally, Crisostomi et al. (2023) study the conditions under which two latent spaces can be merged into a unified one. Moschella et al. (2023) construct a unified space shared by different NNs, enabling zero-shot stitching of independently trained models across different modalities (Norelli et al., 2023). More recently, Cannistraci et al. (2024) enable model stitching without explicit assumptions about the transformation class connecting the latent manifold embeddings, or with only partial correspondence between latent spaces (Cannistraci et al., 2023). Finally, Lähner & Moeller (2024); Maiorca et al. (2024) demonstrate that representations learned by distinct NNs can be aligned using simple transformations.
Architectural Efficiency
While large-scale models with billions or even trillions of parameters continue to achieve state-of-the-art performance, their growth comes with trade-offs, such as slower inference times and significantly higher computational costs. Improving the efficiency of Deep Neural Network (DNN) has been a long-standing area of research. For instance, Veit et al. (2016) shows that removing residual blocks from deep Convolutional Neural Networks (CNNs) only marginally impacts performance, which inspired approaches to reduce inference time by dynamically deciding which layers to execute based on the input (Wu et al., 2018; Veit & Belongie, 2018). Additionally, various techniques to enhance efficiency have emerged, such as early exiting and model pruning. Early exit strategies, which introduce intermediate output layers at different stages of the network, have been shown to reduce inference time (Xin et al., 2020; Zhou et al., 2020; Yu et al., 2022; Tang et al., 2023). However, these approaches require the training of intermediate classifiers to enable exits at predefined layers. Alternatively, model pruning reduces computational load by either removing individual weights based on specific criteria, such as gradient information (Ma et al., 2023), entropy (Liao et al., 2023), or second-order information (Singh & Alistarh, 2020), or by eliminating larger structural components, like channels or residual blocks in ResNets (Bai et al., 2023; Wang & Wu, 2023) and self-attention layers in Transformers (Zhang & He, 2020; Sajjad et al., 2023; Venkataramanan et al., 2024; Zhang et al., 2024). Although effective, these approaches require training the model from scratch and, in the best case, fine-tuning. However, Bai et al. (2023) shows that for CNNs, this additional training step can sometimes be avoided.
Unlike other methods, TOAST leverages intra-network similarities to reduce vision transformers complexity without the need for additional training steps while maintaining competitive performance.
3 Blocks Approximation
The central idea of our approach is that it is possible to leverage representation similarities within transformer-based architectures to replace entire attention blocks with simpler transformations. In this work, a block refers to a sequence of layers including multi-head self-attention, normalization, and feed-forward layers, that function together as a cohesive unit. By replacing these blocks with simpler transformations, we can reduce the computational complexity of the network while maintaining its core functionality.
Approximating Transformer Blocks
Given two blocks and , our goal is to replace the intermediate blocks with a single, lightweight transformation that maps the output of block directly to an approximation of the output of block . This approach allows us to skip the computation of blocks , effectively reducing the overall computational costs. This approximation can be repeated for multiple, non-overlapping blocks, i.e., blocks and with . An overview of the method is provided in Figure˜1.
Let and represent the output representations from block and respectively, for the data points in , sampled uniformly at random from the full training dataset . Our objective is to find a transformation such that:
In this work, we consider to be the identity or a linear transformation . We can compute the linear transformation by minimizing the squared error between the transformed output and the actual :
This optimization problem allows for a closed-form solution that efficiently computes the optimal transformation . The solution bypasses the computation of all layers between any two blocks and , replacing them with . This approximation reduces computational complexity while minimally affecting internal representations, as illustrated in Figures˜9, 7, 10, 11 and 8, and preserves compatibility with downstream classifiers, achieving significant compression as shown in Tables˜1, 2, 3, 7, 8, 9 and 10.
Patterns of Similarity between Transformer Blocks
Inspired by existing results Kornblith et al. (2019); Nguyen et al. (2020), and in particular Venkataramanan et al. (2024), which show that multi-head attention modules exhibit similarity in learned representations, we investigate whether pre-trained foundation models contain entire blocks that produce highly similar representations. To investigate this hypothesis, we adopt the CKA, where a high score between the outputs of two blocks indicates that their respective representations are highly similar. This, in turn, suggests that the intermediate blocks contribute minimally to transforming the representation, highlighting potential redundancy in the model.
After centering the representations, i.e., and , by subtracting their respective means, the CKA is computed as:
Here, denotes the Frobenius norm. A CKA value closer to 1 indicates a higher degree of similarity between the representations of the two blocks.
By systematically evaluating the CKA across pairs of blocks , it is possible to distinguish patterns of similarities between different blocks. These patterns indicate opportunities for targeted complexity reduction, enabling the approximation of redundant blocks without substantially altering internal representations or downstream task performance.
4 Experiments
In this section, we first analyze the similarities between different transformer blocks to motivate their approximation using simple transformations. We then present comprehensive results on image classification across various models and datasets to demonstrate the effectiveness and efficiency of the proposed method. Beyond these core results, we further study the robustness of TOAST through ablations on the number of samples required for approximation and the choice of translator architecture. Overall, our findings show that TOAST achieves strong performance while producing lighter and faster models. Due to space constraints, additional results on zero-shot image classification, as well as further qualitative and quantitative analyses, are provided in the Appendix (Sections˜A.2.3 and A.2.2).
4.1 Latent analysis
In this section we investigate similarities in the latent representations of DiNO-B and DEiT-S on five datasets: CIFAR-10, CIFAR-100, MNIST, F-MNIST, and ImageNet1k. We compute the CKA similarity metric averaged over all tokens from a small subset of 100 training samples, which is sufficient to reveal block-level similarity patterns while remaining computationally efficient. Additional results with other pretrained vision transformers (ViT-T, ViT-S, DiNO-S, ViT-B) are provided in Section˜A.2.1, showing consistent patterns for each model across different datasets.
Do vision transformer models exhibit block-wise similarity patterns?
The results in Figure˜2 reveal that while the similarity patterns differ across models, they remain largely consistent for the same model across different datasets. This suggests that the similarity structure between computational blocks is predominantly influenced by the model itself. Although the general similarity pattern remains the same, the differences in CKA values become more pronounced (i.e., the block structure becomes more evident) as the complexity of the dataset increases (e.g., from MNIST to ImageNet1k). These finding aligns with observations from Nguyen et al. (2020), where DNN trained from scratch exhibit a distinctive "block structure" in their representations, which is linked to model overparameterization. Our results extend this observation to vision pre-trained foundation models, showing that such a structure is primarily an intrinsic property of the model. Moreover, these consistent block-wise patterns indicate potential targets for approximation, suggesting that entire blocks may be replaced with simpler transformations without substantially altering the model’s internal representations.
How does TOAST affect latent representations?
We next analyze the impact of the proposed transformations on the final block’s latent representations, which are used for downstream classification. We approximate these blocks using a shared linear transformation applied across all tokens, estimated on a subset of 500 training samples. For consistency, we use the same models and datasets as in Figure˜2.
Original
TOAST
DiNO-B
DEiT-S
To quantify the effect of the approximation, we compute the CKA similarity between the final block representations of the original and the TOAST-approximated model for each block using its preceding block as input. As shown in Figure˜4, the model-specific similarity patterns re-emerge after approximation. The plots highlight more specific trends. Approximating blocks is easier on simpler tasks (e.g., image classification on MNIST or F-MNIST), yielding representations that closely match the originals, whereas on more complex datasets (e.g., ImageNet1k or CIFAR-100), the approximated representations deviate more from the original ones. Furthermore, the final blocks of DEiT-S exhibit high similarity, suggesting that approximating these layers preserves the final representations, while earlier blocks remain more critical.
To provide a more intuitive view, Figure˜4 visualizes the final-layer representations using Principal Component Analysis (PCA). We compare the original representations with those obtained after approximating the final block (10 11) using TOAST on F-MNIST, with colors indicating the 10 classes. The visualization confirms that approximating the final block of DiNO-B results in noticeable deviations from the original representations, whereas for DEiT-S the approximated representations remain highly similar. These observations align with the CKA analysis in Figure˜4, highlighting that the effect of block approximation depends strongly on the model and its internal block structure. Additional results across other models and datasets are provided in Section˜A.2.1.
| Identity Translator | Linear Translator | |||||||
| Approx. | Params. | Accuracy % | GFLOPS | imgs/s | Accuracy % | GFLOPS | imgs/s | |
| 2 4 | -3.25M | |||||||
| 9 11 | -3.25M | |||||||
| 0 1 | -1.62M | |||||||
| 10 11 | -1.62M | |||||||
| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
Can entire transformer blocks be approximated without losing accuracy?
Initial results, reported in Table˜1, support the qualitative analysis and empirically demonstrate that entire vision transformer blocks can be effectively approximated using simpler transformations (e.g., linear projections or, in some cases, the identity function). Such approximations reduce both the number of parameters and Giga Floating-Point Operations (GFLOPs), thereby improving throughput (images per second), while incurring only a slight to negligible decrease in downstream task performance. For instance, consistent with our earlier analysis, we find that approximating the final block of DEiT-S when using ImageNet1k (e.g., approximating blocks 10 11 or 9 11 with a linear transformation) yields modest performance drops of -0.09% and -3.38%, respectively, while providing substantial efficiency gains. Importantly, we also show that even the identity transformation, achieves competitive results, with accuracy drops as small as -0.24% and -5.44%, respectively. However, the choice of translator naturally depends on the efficiency-accuracy trade-off: linear translation guarantee in general most reliable accuracy–efficiency balance, whereas the identity yields the leanest training-free approximation when maximum simplicity is required. Further methodological details and the full evaluation are presented in Section˜A.1 and Section˜4.2 respectively, while details on the efficiency metrics and additional analysis on those are in Sections˜A.1.5 and A.2.5, respectively.
4.2 Image Classification Performance
We evaluate TOAST on image classification tasks using pretrained models of varying sizes (ViT-L, DiNO-B, and DEiT-S) and two benchmark datasets (CIFAR-100F and ImageNet1k). Additional results with a broader set of models (ViT-T, ViT-S, ViT-B, ViT-L, DiNO-S, DiNO-B, DEiT-S) and datasets (MNIST, F-MNIST, CIFAR-10, CIFAR-100C) are provided in Section˜A.2.2. While in Section˜A.2.6, we complement the quantitative evaluations with qualitative analyses of misclassifications after block approximation, providing further insight into model behavior under TOAST. Additional implementation details, including model and dataset specifications, computational resources, and software tools, are provided in Tables˜5 and 6, and Sections˜A.1.4, A.1.5 and A.1.3.
| Identity | Linear | |||||||
| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
DEiT-S |
3 4, 9 11 | -4.88M | 3.74 | 7751.4 | 3.80 | 7718.4 | ||
| 3 4, 9 10 | -3.25M | 4.15 | 7210.9 | 4.21 | 7188.4 | |||
| 2 3 | -1.62M | 4.56 | 6754.2 | 4.59 | 6736.7 | |||
| 10 11 | -1.62M | 4.56 | 6752.6 | 4.59 | 6740.5 | |||
| original | 21.81M | 4.97 | 6349.2 | 4.97 | 6325.6 | |||
DiNO-B |
0 4 | -26.00M | 16.32 | 3230.9 | 16.47 | 3227.7 | ||
| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2947.0 | 18.80 | 2929.6 | |||
| 0 1, 2 3 | -13.00M | 20.37 | 2703.9 | 20.67 | 2681.2 | |||
| 0 1 | -6.50M | 22.39 | 2506.6 | 22.54 | 2487.0 | |||
| 5 6 | -6.50M | 22.39 | 2503.1 | 22.54 | 2490.6 | |||
| original | 86.58M | 24.42 | 2321.3 | 24.42 | 2316.5 | |||
ViT-L |
2 4, 18 23 | -80.83M | 45.05 | 1654.9 | 45.47 | 1652.8 | ||
| 17 23 | -69.28M | 47.70 | 1572.4 | 47.90 | 1567.0 | |||
| 3 4, 19 23 | -57.74M | 50.34 | 1509.9 | 50.75 | 1499.5 | |||
| 3 4, 20 23 | -46.19M | 52.98 | 1440.4 | 53.39 | 1436.8 | |||
| 3 4, 21 23 | -34.64M | 55.62 | 1377.2 | 56.03 | 1345.6 | |||
| 7 8, 15 16 | -23.09M | 58.26 | 1318.2 | 58.67 | 1316.4 | |||
| 16 17, 22 23 | -23.09M | 58.26 | 1318.8 | 58.67 | 1312.3 | |||
| 3 4 | -11.55M | 60.90 | 1269.2 | 61.11 | 1270.0 | |||
| 22 23 | -11.55M | 60.90 | 1267.5 | 61.11 | 1270.9 | |||
| original | 304.35M | 63.54 | 1219.8 | 63.54 | 1225.2 | |||
Block approximations in TOAST are calculated via a shared linear, or identity, transformation applied across all tokens and are estimated using a subset of 500 training samples. A linear classifier is then trained on top of the frozen backbone with the Adam optimizer (learning rate ), batch size 256, for 5 epochs, over 3 different seeds. This setup simulates a realistic scenario where a pretrained feature extractor is adapted to a new dataset unseen during pretraining. However, to assess the robustness of our method, we also report the results using the original classification heads (Section˜A.2.4), which confirm the consistency of our findings.
Are TOAST results competitive?
As shown in Tables˜3 and 2, TOAST consistently reduces model size and GFLOPs while maintaining, and in some cases improving, image classification accuracy. This aligns with our representational analyses in Section˜4.1: for instance, approximating the final block of DEiT-S produces latent representations nearly identical to the original (Figures˜4 and 4), making it an ideal candidate for approximation. Even when multiple consecutive blocks are approximated (e.g., 911), models maintain performance comparable to or exceeding the original while significantly reducing parameters. This demonstrates that a simple linear transformation, or even the identity in certain cases, is sufficient to capture the functionality of full transformer blocks without additional training, provided the transformation is shared across all tokens. Additionally, efficiency gains are notable: throughput (imgs/s) increases while GFLOPs decreases, highlighting practical benefits for deployment, as also shown in Section˜A.2.5. Additional results across other models (DiNO-B, ViT-L) and datasets confirm that TOAST generalizes across architectures and scales (Section˜A.2.2). Finally, while approximations are easier for simpler datasets (e.g., CIFAR-100F), TOAST still achieves meaningful compression with minimal accuracy loss on complex datasets like ImageNet1k.
| Identity | Linear | |||||||
| Approx. | Params. | Accuracy % | GFLOPs | imgs/s | Accuracy % | GFLOPs | imgs/s | |
DEiT-S |
3 4, 9 11 | -4.88M | 3.74 | 7755.1 | 3.80 | 7713.7 | ||
| 9 11 | -3.25M | 4.15 | 7226.6 | 4.18 | 6791.7 | |||
| 8 9 | -1.62M | 4.56 | 6755.2 | 4.59 | 6739.9 | |||
| 9 10 | -1.62M | 4.56 | 6692.1 | 4.59 | 6741.3 | |||
| original | 21.81M | 4.97 | 6349.0 | 4.97 | 6249.4 | |||
DiNO-B |
0 4 | -26.00M | 16.32 | 3233.8 | 16.47 | 3204.9 | ||
| 0 1, 2 3, 4 5 | -19.50M | 18.34 | 2943.1 | 18.80 | 2922.6 | |||
| 0 1, 2 3 | -13.00M | 20.37 | 2705.6 | 20.67 | 2690.1 | |||
| 0 1 | -6.50M | 22.39 | 2492.8 | 22.54 | 2493.8 | |||
| 2 3 | -6.50M | 22.39 | 2484.2 | 22.54 | 2484.6 | |||
| original | 86.58M | 24.42 | 2315.5 | 24.42 | 2317.3 | |||
ViT-L |
2 4, 18 23 | -80.83M | 45.05 | 1655.7 | 45.47 | 1649.6 | ||
| 17 23 | -69.28M | 47.69 | 1578.8 | 47.90 | 1552.1 | |||
| 3 4, 19 23 | -57.74M | 50.34 | 1503.6 | 50.75 | 1497.4 | |||
| 3 4, 20 23 | -46.19M | 52.98 | 1445.2 | 53.39 | 1431.0 | |||
| 20 23 | -34.64M | 55.62 | 1381.2 | 55.82 | 1372.6 | |||
| 3 4, 21 23 | -34.64M | 55.62 | 1376.7 | 56.03 | 1372.7 | |||
| 20 22 | -23.09M | 58.26 | 1317.5 | 58.47 | 1314.6 | |||
| 3 4, 21 22 | -23.09M | 58.26 | 1315.8 | 58.67 | 1317.6 | |||
| 20 21 | -11.55M | 60.90 | 1268.5 | 61.11 | 1266.7 | |||
| 21 22 | -11.55M | 60.90 | 1270.7 | 61.11 | 1269.2 | |||
| original | 304.35M | 63.54 | 1223.1 | 63.54 | 1224.3 | |||
Additional results across models and datasets are provided in Tables˜7, 8, 9 and 10. To assess scalability, we applied TOAST to ViT-L. Approximating selected blocks, e.g., 17 23, reduces the parameter count by 69.3M, lowers GFLOPs from 63.54 to 47.79, and increases throughput from 1223.1 to 1578.8 imgs/s, while incurring a minimal accuracy drop of 1.16%. This shows TOAST’s utility in balancing substantial computational savings with a modest performance trade-off, even in large models.
Are 500 training samples enough?
We study the sensitivity of block approximation to the number of training samples using DiNO-B and DEiT-S on ImageNet1k. As shown in Figure˜5, performance typically plateaus quickly: 500 samples are sufficient to obtain stable and reliable approximations. Increasing the sample count beyond this threshold provides only marginal gains, while substantially fewer samples lead to noticeable degradation. Interestingly, when the representational spaces of consecutive blocks are already highly aligned, even as few as 10 or 50 samples suffice to achieve competitive approximations. Conversely, for blocks that are harder to approximate, such as the early layers of DEiT-S (e.g., 01), even 4000 samples are insufficient to estimate a linear transformation that maintains competitive performance. We highlight that these results are obtained on ImageNet1k, which contains 1000 classes. The 500 samples represent only a small subset of the class space, yet reliable approximations are still achieved. This indicates that TOAST primarily captures the block-level structure of representations rather than requiring exhaustive coverage of all classes. Consequently, TOAST could be practical also in scenarios where a large labeled datasets is limited.
What if a more complex transformation is used?
We evaluate whether deeper approximators improve downstream task performance. Specifically, we compare TOAST (Identity and Linear) to MultiLayer Perceptron (MLP) and Residual MLP, trained for 300 steps with Adam (learning rate ). These more complex transformation, as for Identity and Linear, are applied across all tokens, and estimated using a subset of 500 training samples. Results in Table˜4 show a consistent trend for ViT-L on both ImageNet1k and CIFAR-100F: the linear transformation provides the most reliable trade-off across datasets. On CIFAR-100F, linear often achieves the best or near-best accuracy (e.g., 2122: vs. for Res-MLP and for MLP), while remaining training-free, thus more efficient. On ImageNet1k, the gap becomes even clearer: for the same blocks linear reaches , while Res-MLP and MLP reach and , respectively. Additionally, also Linear obtain competitive results. TOAST operates in closed form, requires no optimization, and consistently achieves strong efficiency–accuracy trade-offs. These findings confirm that a simple linear transformation is sufficient to approximate transformer blocks in most settings, with deeper translators offering little benefit despite their higher cost.
| Accuracy | ||||||
| Approx. | Params. | Identity | Linear | MLP | Res-MLP | |
CIFAR-100F |
3 4, 20 23 | -46.19M | ||||
| 17 23 | -69.28M | |||||
| 3 4, 19 23 | -57.74M | |||||
| 20 23 | -34.64M | |||||
| 20 22 | -23.09M | |||||
| 3 4, 21 22 | -23.09M | |||||
| 20 21 | -11.55M | |||||
| 21 22 | -11.55M | |||||
| original | 304.35M | |||||
ImageNet1k |
3 4, 20 23 | -46.19M | ||||
| 17 23 | -69.28M | |||||
| 3 4, 19 23 | -57.74M | |||||
| 20 23 | -34.64M | |||||
| 20 22 | -23.09M | |||||
| 3 4, 21 22 | -23.09M | |||||
| 20 21 | -11.55M | |||||
| 21 22 | -11.55M | |||||
| original | 304.35M | |||||
5 Limitations and future work
While TOAST efficiently approximates transformer blocks, our current investigation has primarily focused on vision transformer architectures and their application to classification tasks. Future research will explore the applicability of TOAST to other modalities (e.g., text) and to diverse downstream tasks (e.g., image reconstruction). Such an expansion will be crucial for testing the universality of the observed block-similarity phenomena and assessing TOAST’s adaptability. Furthermore, we aim to expand the analysis of these block-level similarities. This involves investigating redundancies at finer granularities, such as within individual attention heads or feed-forward layers, and consistently and automatically identifying which blocks to approximate, resulting in minimal degradation of downstream task performance. Another important direction is to examine more in detail how these similarity patterns arise depending on token type (e.g., CLS token versus mean token). The analysis could lead to even more refined and context-aware approximation strategies, further enhancing model efficiency.
6 Conclusion
In this work, we first analyze the emergence of consistent block-wise representation similarities within pretrained foundation models and then propose a method to leverage these similarities to obtain smaller and more efficient yet performant models. To this end, we propose Transformer Optimization using Adaptive and Simple Transformations (TOAST), a novel method for easily approximate entire transformer blocks using a simple transformation, without requiring additional training or fine-tuning. Our extensive empirical evaluations across multiple pretrained vision models and datasets validate that TOAST significantly reduces model parameters while maintaining, and sometimes even improving, downstream task performance. Furthermore, TOAST’s straightforward linear approach often achieves better results than existing strategies like block skipping, and can be more effective than complex, trained approximations. TOAST thus offers a practical and efficient method for streamlining foundation models, making them more computationally accessible, and towards deployment in resource-constrained scenarios such as on-device settings.
Acknowledgments
The authors gratefully acknowledge Luca Moschella for the insightful discussions. TS and EP are supported by the grant #2021-911 of the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” of the ETH Domain (Swiss Federal Institutes of Technology). EP is also supported by a fellowship from the ETH AI Center. This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI).
Appendix A Appendix
A.1 Implementation details
This section details the experiments conducted in Section˜4, providing information to reproduce them. Additionally, we provide the code as a zip file in the supplementary material.
A.1.1 Models and Datasets
Table˜5 contains the full list of the pretrained models, while Table˜6 contains dataset information.
| Modality | HuggingFace Model Name | Alias | Enc. Dim |
| Vision | WinKawaks/vit-tiny-patch16-224 | ViT-T (Dosovitskiy et al., 2021) | 192 |
| WinKawaks/vit-small-patch16-224 | ViT-S (Dosovitskiy et al., 2021) | 384 | |
| facebook/dinov2-small | DiNO-S (Oquab et al., 2023) | 384 | |
| facebook/deit-small-patch16-224 | DEiT-S (Touvron et al., 2020) | 384 | |
| google/vit-base-patch16-224 | ViT-B (Dosovitskiy et al., 2021) | 768 | |
| facebook/dinov2-base | DiNO-B (Oquab et al., 2023) | 768 | |
| laion/CLIP-ViT-B-16-laion2B-s34B-b88K | OpenCLIP-ViT-B (Zhai et al., 2019) | 768 | |
| google/vit-large-patch16-224 | ViT-L (Dosovitskiy et al., 2021) | 1024 |
| Modality | Name | Alias | Classes |
| Vision | MNIST (Deng, 2012) | MNIST | 10 |
| Fashion-MNIST (Xiao et al., 2017) | F-MNIST | 10 | |
| CIFAR-10 (Krizhevsky et al., 2009) | CIFAR-10 | 10 | |
| CIFAR-100 (coarse) (Krizhevsky et al., 2009) | CIFAR-100C | 20 | |
| CIFAR-100 (fine) (Krizhevsky et al., 2009) | CIFAR-100F | 100 | |
| Imagenet-1k (Russakovsky et al., 2015) | ImageNet1k | 1000 |
A.1.2 Approximators
The first implementation, referred to as the Res-MLP, is composed of two normalization layers and a feedforward submodule. The first layer normalization processes the input tensor, followed by a feedforward submodule comprising a linear transformation, a SiLU activation, a dropout layer, and a final linear transformation. The output of the feedforward submodule is added to the normalized input via a residual connection. This sum is then passed through the second normalization layer to produce the final output. The second implementation, referred to as the MLP, is a simplified MLP that employs a sequential architecture with a first linear transformation that reduces the input dimensionality to half of the target dimension, followed by a GELU activation function for smooth non-linearity, and a final linear transformation that restores the reduced features to match the target dimensionality. Refer to LABEL:lst:mlp1 and LABEL:lst:mlp2 for the code snippet of the two translators.
A.1.3 Tools & technologies
All the experiments presented in this work employ the following tools:
PyTorch Lightning, to ensure reproducible results while also getting a clean and modular codebase;
NN-Template GrokAI (2021), to easily bootstrap the project and enforce best practices;
Transformers by HuggingFace, to get ready-to-use transformers for both text and images;
Datasets by HuggingFace, to access most of the datasets;
DVC (Kuprieiev et al., 2022), for data versioning;
fvcore analysis library , for calculating GFLOPs;
A.1.4 Computational resources
Experiments involving larger models, specifically DiNO-B, OpenCLIP-ViT-B, and ViT-L, were conducted on an NVIDIA H100 GPU equipped with 93 GB of memory. All the other experiments utilized an NVIDIA GeForce RTX 5090 GPU with 31 GB of memory.
A.1.5 Efficiency metrics
We evaluated model efficiency using two primary metrics. GFLOPs were used to measure the hardware-independent theoretical complexity of a single forward pass, calculated using the fvcore analysis library. Throughput, measured in samples per second, was used to quantify the practical, hardware-dependent inference speed. This was benchmarked by averaging the wall-clock time over numerous iterations on a single NVIDIA H100 GPU with a consistent batch size of 256.
A.2 Additional Experiments
This section presents supplementary experiments to extend those detailed in Section˜4.
A.2.1 Latent Analysis
This section extend the analysis conducted in Section˜4.1, to analyze block-wise internal similarities, to additional models of different dimensionality: ViT-T, ViT-S, ViT-B and DiNO-S. Additionally, we provide visualization using PCA for DiNO-S, DEiT-S, ViT-S, with different datasets and approximating both early and late blocks (see Figures˜10, 11, 9, 7 and 8).
A.2.2 Image Classification
This section presents additional experiments that complement and extend those detailed in Section˜4.2. Datasets and models are the ones detailed in Tables˜6 and 5.
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F | ImageNet1k |
| 1 5 | 15.31M | ||||||
| 2 5 | 16.94M | ||||||
| 7 10 | 16.94M | ||||||
| 1 3 | 18.56M | ||||||
| 2 4 | 18.56M | ||||||
| 3 5 | 18.56M | ||||||
| 1 2, 3 4 | 18.56M | ||||||
| 1 2, 4 5 | 18.56M | ||||||
| 0 1 | 20.43M | ||||||
| 1 2 | 20.43M | ||||||
| 2 3 | 20.43M | ||||||
| 3 4 | 20.43M | ||||||
| 4 5 | 20.43M | ||||||
| 5 6 | 20.43M | ||||||
| 6 7 | 20.43M | ||||||
| 7 8 | 20.43M | ||||||
| 8 9 | 20.43M | ||||||
| 9 10 | 20.43M | ||||||
| 10 11 | 20.43M | ||||||
| original | 22.06M |
| Accuracy | ||||||
| Approx. | Params. | MNIST | F-MNIST | CIFAR-10 | CIFAR-100C | CIFAR-100F |
| 1 5 | -25.99M | |||||
| 2 5 | -19.49M | |||||
| 1 3 | -13M | |||||
| 3 5 | -13M | |||||
| 2 4 | -13M | |||||
| 8 10 | -13M | |||||
| 9 11 | -13M | |||||
| 3 4 | -6.5M | |||||
| 4 5 | 6.5M | |||||
| 0 1 | -6.5M | |||||
| 1 2 | -6.5M | |||||
| 2 3 | -6.5M | |||||
| - | 86.39M | |||||
A.2.3 Zero-shot Image Classification
To further assess the effectiveness of our approach, we evaluate TOAST in a zero-shot image classification setting. This evaluation utilizes the OpenCLIP-ViT-B model (Radford et al., 2021), which was pretrained on LAION-2B Schuhmann et al. (2022), with ImageNet1k serving as the downstream evaluation dataset.
| Params. | Approx. | Accuracy | |
| -6.49M | 0 1 | 57.93 | - |
| 1 2 | 64.20 | - | |
| 2 3 | 66.35 | -5.51% | |
| 3 4 | 64.65 | - | |
| 4 5 | 64.86 | - | |
| 5 6 | 58.05 | - | |
| 6 7 | 61.56 | - | |
| 7 8 | 58.53 | - | |
| 8 9 | 52.32 | - | |
| 9 10 | 59.21 | - | |
| 10 11 | 22.64 | - | |
| 149.07M | original | 70.21 | – |
The analysis is conducted only on the base version, as larger versions (e.g., OpenCLIP-ViT-L or OpenCLIP-ViT-H) contain too many parameters and are thus beyond the scope of this paper. As in previous experiments, the model remains frozen, and block approximations are computed using a shared linear transformation applied across all tokens, based on a subset of 3,000 training samples. Importantly, we apply these approximations only to the vision encoder, leaving the text encoder unchanged. We follow the standard ImageNet1k prompt templates. The results in Table˜11 lead to the conclusion that the impact on zero-shot accuracy is highly dependent on the targeted block’s position. The choice of which blocks to approximate is therefore crucial. For instance, approximating an early block (e.g., 1 2 or 2 3) results in a modest accuracy drop (e.g., 5.51%), yielding a competitive model with fewer parameters. In contrast, approximating the final block (i.e., 10 11) causes a catastrophic performance collapse of 67.75%. Meaning that, for OpenCLIP-ViT-B, later layers in the vision encoder appear to capture uniquely critical information for zero-shot generalization that cannot be effectively replicated by earlier ones. To the best of our knowledge, this work is the first to investigate training-free model size reduction in this challenging setting.
A.2.4 Evaluation with original classification heads
| Encoder | Approximation | Accuracy |
DEiT-S |
78.87 | |
| 78.72 | ||
| 78.39 | ||
| 75.29 | ||
| 74.36 | ||
| 72.34 | ||
| 69.30 | ||
| 66.12 | ||
| original | 79.66 | |
ViT-S |
78.25 | |
| 74.50 | ||
| 71.23 | ||
| 70.46 | ||
| 66.18 | ||
| 59.69 | ||
| 58.36 | ||
| 23.85 | ||
| original | 79.86 |
As mentioned in the main paper, our primary evaluation involves training a new linear classifier on top of the frozen model backbone to simulate a realistic transfer learning scenario. However, the original papers for DEiT-S Touvron et al. (2021) and ViT-S Beyer et al. (2022) report performance using the classification head that was part of the original pre-training.
To confirm that our conclusions are robust and not an artifact of our evaluation protocol, we conducted an additional set of experiments using the official, pre-trained classification heads from the original model checkpoints. In this setup, we do not train a new classifier; we simply evaluate the accuracy of the frozen, approximated models using their original heads.
The results, presented in Table˜12, are fully consistent with the main conclusions of our paper. They confirm that our block approximation method provides a favorable accuracy-efficiency trade-off, even when evaluated with the original model heads. The relative drop in accuracy when approximating different layers follows the same patterns observed in our primary experiments, reinforcing the validity of our approach.
A.2.5 Computational efficiency vs. accuracy
To quantify the effectiveness of different approximation methods, we analyze the trade-off between downstream accuracy and computational cost. Figure˜12 presents this analysis on a DiNO-B model using both CIFAR-100F and ImageNet1k against three standard efficiency metrics: parameter count, GFLOPs, and inference throughput. Across all metrics, the proposed linear translator (green) establishes a more favorable Pareto frontier compared to the baseline identity-based approach (blue). This indicates that for any given efficiency budget (e.g., a specific GFLOPs target), the linear translator consistently yields a model with higher accuracy.
A.2.6 Analysis of misclassifications
In this section, we examine changes in per-class accuracy and misclassification patterns. As shown in Figure˜13, models behave differently at block approximations. DiNO-S remains remarkably stable across blocks and classes, with the only degradation appearing for classes dog (when approximating blocks 10 or 11) and deer (for block 10 approximation). ViT-S shows a similar drop for class dog on its final block. Instead, the most noticeable hit occurs for class cat when the earlier blocks are approximated. For DEiT-S, several mid‑to‑late block approximations improve accuracy for various classes, whereas the very first block causes a clear relative decline in nearly every class. These observations suggest strategies like preferring late‑block approximation for DEiT-S, or reserving extra samples for the linear transformation in order to recover the accuracy of difficult classes for the model.
In order to further investigate how the predictions change while approximating blocks, we plot the difference in the normalized confusion matrix before and after the approximation. In Figure˜14, we show the delta confusion matrix for DEiT-S on CIFAR-100C. Also, here we can see how approximating the very first block makes the model puzzling and lose per-class accuracy (i.e., negative delta along the diagonal). On the other hand, approximating the last block acts as a regularizer, resulting in an overall gain in the per-class accuracy and, as a consequence, fewer misclassifications (negative deltas off-diagonal). This supports results shown in Figures˜13 and 3.
Additionally, Figure˜15 shows representative CIFAR-10 images that become misclassified after approximating a block of ViT-S. The patterns we observe mirror the trends in Figures˜13 and 14: when approximating earlier blocks, we observe many images belonging to class cat to be misclassified. Instead, when approximating later blocks, we observe images of the class dog to be misclassified. Together, these qualitative examples show that understanding these block-specific vulnerabilities allows us to steer the approximation procedure, informing choices about which blocks to approximate based on the observed impact on the final model’s class-wise performance.